14. What is the expected price of a stock which pays a $1 dividend next year, has a required return of 5% and a dividend growth of 1%?

What is the expected total return of the stock?

What is the dividend yield of the stock?

What is the capital gains return of the stock?
Answer #1
  • From the above problem,

dividend(d1) = $1

required rate of return or cost of equity capital (r) = 5% or 0.05

dividend growth (g) = 1% or 0.01

To calculate the expected price of the stock we will use dividend growth model:

Price (P) = d1/(r-g) = $1 / (.05-.01) = $25.

  • Total return = (Return from increase in price of the stock + dividend yield)

Total return = (0.05 * $25 + $1) = $2.25 (in one year.)

  • Dividend Yield = (Dividend per share / Current market price)*100

dividend yield = ($1 / $25)*100 = 4%

  • Capital gain return = $2.25/$25 = 0.09 = 9%